### v1

# Concept – An Evaluation Protocol on Conversation Recommender Systems with System- and User-centric Factors

Abstract
The conversational recommendation system (CRS) has been criticized regarding its user experience in real-world scenarios, despite recent significant progress achieved in academia. Existing evaluation protocols for CRS may prioritize system-centric factors such as effectiveness and fluency in conversation while neglecting user-centric aspects. Thus, we propose a new and inclusive evaluation protocol, Concept, which integrates both system- and user-centric factors. We conceptualise three key characteristics in representing such factors and further divide them into six primary abilities. To implement Concept, we adopt a LLM-based user simulator and evaluator with scoring rubrics that are tailored for each primary ability. Our protocol, Concept, serves a dual purpose. First, it provides an overview of the pros and cons in current CRS models. Second, it pinpoints the problem of low usability in the "omnipotent" ChatGPT and offers a comprehensive reference guide for evaluating CRS, thereby setting the foundation for CRS improvement.
Concept – An Evaluation Protocol on Conversation Recommender Systems with System- and User-centric Factors
[…]
1 Introduction
The synergies between the conversation interface and recommendation system have given rise to a groundbreaking paradigm known as the Conversational Recommendation System (CRS) Sun and Zhang (2018); Gao et al. (2021). It acts as a cooperative agent that engages in a natural language conversation with users and provides recommendations. Despite the great success achieved, CRS has been criticized regarding its user experience in real-world scenarios, lacking practical usability Jannach and Manzoor (2020); Jannach et al. (2021). This is partly due to the fact that current system-centric evaluation protocols tend to prioritize assessing the characteristics of the CRS system per se, such as response diversity and fluency Ghazvininejad et al. (2018); Wang et al. (2022a), as well as the recommendation effectiveness and efficiency Wang et al. (2023d); Jin et al. (2019); Wärnestål (2005). Such protocols often overlook user-centric factors, which gauge how users engage with and perceive the social capabilities of the CRS. For instance, a CRS model may provide accurate recommendations and fluent conversations but at the same spread dishonest information. This can be misleading to users resulting in an unsatisfactory user experience. Therefore, it is imperative to consider incorporating both system- and user-centric factors in the evaluation protocol to develop a more user-friendly CRS system.
Figure 1: Concept integrates both system- and user-centric factors into three characteristics based on the previous taxonomy on human-AI interactions. Such characteristics are further divided into six primary abilities to enhance the inclusiveness in evaluations.
To resolve this issue, we trace back the taxonomy that examines how the factors of conversational AI impact the user experience in human-AI interactions Chaves and Gerosa (2021); Reeves and Nass (1996), and tailor a particular evaluation protocol for CRS. We introduce Concept, an COmpreheNsive CRS Evaluation ProTocol. As framed in Figure 1, Concept considers both system- and user-centric factors and conceptualizes them into three characteristics, which are further divided into six primary abilities. Such hierarchical factors are taken as inclusive and fine-grained evaluations. In addition, we present a practical implementation of Concept utilizing an LLM-based user simulator and evaluator, together with automated computational metrics. Specifically, the simulator mimics human social cognition and interacts with CRS to generate conversation data. Then, the evaluator assigns scores based on ability-specific scoring rubrics. This enables Concept to conduct controlled and inclusive evaluations.
By applying Concept, we can evaluate and analyze the strengths, weaknesses, and potential risks of off-the-shelf CRS models [^1] We conduct evaluations using both humans and LLM, and found highly correlated results of the two.. A total of 6720 conversation data is recorded to collect the interactions between off-the-shelf CRS and simulated users who demonstrate different personas and preferences. Existing works agree on the insurmountable performance achieved by ChatGPT-based CRS. However, current CRS models, even enhanced by "omnipotent" ChatGPT, still encounter numerous challenges, falling short of practical usability in particular. 1) They struggle to express genuine responses without hallucination or deceit, often introducing non-existent items into conversations and present to users. 2) They lack of self-awareness of its identity, facing difficulties in producing both persuasive and honest explanations, which is prominent in the ChatGPT-based CRS model, where explanations are highly convincing but frequently contain illusory details, misleading users in believing that these items align with their preferences. 3) They encounter issues in offering reliable recommendations as they are sensitive to contextual nuances. Even slight alterations in user wording may result in entirely different recommendations. 4) They lack proficiency in catering to diverse users without prior coordination, failing to dynamically adjust their behavior to align with each user’s distinct personas. It is observed that the ChatGPT-based CRS model tends to employ deceptive tactics to persuade optimistic users to accept recommendations. This underscores the importance of aligning CRS with human values and advocating its ethical use.
With Concept, we provide an overview for researchers as a reference guidance for evaluating CRS and laying the foundation for CRS enhancement. The main contributions as follows:
  • We pinpoint the fact that making a CRS admirable to users is primarily a social problem, instead of just a technical one. Social attributes are the key to the widespread acceptance of CRSs.
  • We initiate the work on conceptualizing CRS’s characteristics in a comprehensive way, combining both system- and user-centric factors.
  • We propose a new evaluation protocol, called, Concept, which conceptualizes user expectations into six abilities, together with a scoring implementation.
  • We evaluate and analyze the strengths, weaknesses, and potential risks of off-the-shelf CRS models in order to provide a fundamental perspective for people to make a reference for CRS evaluation.
2 Related Work
Our research focuses on the CRS evaluation, with LLM-based methods in particular. Hence, our literature review discusses the same scope of research. Most relevant works are summarized below, and more detailed review is attached in Appendix B.
CRS Evaluation. We reckon that the success of CRS in practice may rely primarily on social characteristics, instead of just technical ones. However, existing evaluation protocols mainly focus on system-centric evaluation aspects, such as lexical diversity and perplexity of responses Ghazvininejad et al. (2018); Chen et al. (2019), or conversational fluency, relevance, and informativeness Wang et al. (2022b, a), or recommendation effectiveness and efficiency Wang et al. (2023d); Jin et al. (2019); Wärnestål (2005) that are often straightforward and function-oriented. These aspects seem unable to reveal the inadequacy of CRS from user-centric perspectives. Hence, the fragmented evaluation protocols are underdeveloped to conduct an inclusive evaluation. Although some efforts have attempted to consider the user-centric characteristics of CRS, they rely on person-to-person conversation analysis and questionnaire interviews, lacking quantitative viewpoints and empirical evidence Jin et al. (2023, 2021); Jannach (2022); Siro et al. (2023) in addition that they might have over-used system-centric characteristics [^2] Refer to Table 8 for better understanding our differences.. In a nutshell, existing protocols may underestimate the problems of evaluating CRS Jannach et al. (2021); Jannach (2023), and our Concept considers both system- and user-centric factors, aiming to provide an inclusive evaluation protocol for CRS.
LLM-based Evaluation. LLM-based evaluation is widely known to be labor-intensive and effective Huang et al. (2023); Afzali et al. (2023); Zeng et al. (2023). Recent studies reach a consensus on emphasizing the importance of providing detailed scoring rubrics in prompts for consistent evaluations to align with human assessments Liu et al. (2023d); Wang et al. (2023d). In line with the above research, Concept also utilizes the LLM for evaluation implementations. Moreover, our simulator is equipped with the Theory of Mind Fischer (2023), which enables the simulator to reflect on its predefined personas.
3 Concept
An inclusive evaluation protocol is essential for understanding the inadequacy of CRS in practice. We present Concept, which consolidates both system- and user-centric factors into three characteristics and six specific abilities in Section 3.1. We outline the evaluation process in Section 4.2.
3.1 Factors Categorization
Interdisciplinary research on conversational AI Chaves and Gerosa (2021); Reeves and Nass (1996); Fogg (2003) groups its characteristics that impact the user experience into three types. Concept then customizes them for CRS in this section, as depicted in Figure 1. For details on the customization, refer to Appendix C.
3.1.1 Recommendation Intelligence
From a system-centric factors, this requires CRS to learn from conversations and evolve toward making recommendations as the conversation advances  Chen et al. (2019); Ma et al. (2020); Zhou et al. (2021). In this regard, our categorization encompasses two primary abilities:
  • Quality. CRS should provide precise recommendations using minimal conversation turns, which are the crucial aspects that influence user satisfaction Siro et al. (2023); Gao et al. (2021). We emphasize the importance of the user acceptance rate, reflecting the practical effectiveness of the recommendations.
    In our experiments, we utilize computational metrics for automatic evaluation, as seen in prior studies Wang et al. (2023d); Zhang et al. (2023); Yu et al. (2023). These metrics encompass $Recall@k(k=1,10,25,50)$, recommendation success rate $(SR@k,k=3,5,10)$, user acceptance rate, and average turns ($AT$) needed to achieve successful recommendations.
  • Reliability. CRS should deliver robust and consistent recommendations that account for contextual nuances. In practical situations, users with similar preferences may express themselves differently. It would diminish the user experience and pose a disruption for critical applications if inconsistent items were recommended for two similar user responses Tran et al. (2021); Oh et al. (2022). In less critical cases, if two recommended items are inconsistent but align with user preferences, they are often viewed as diverse recommendations, creating opportunities for relevant but less popular items Yang et al. (2021).
    To evaluate reliability, we generate sets of user response pairs with similar meanings using ChatGPT paraphrasing. Given a user response pair [$u_{1}$, $u_{2}$], we define four computational metrics: the rate of Consistent Action, indicating whether the CRS constantly provides recommendations based on $u_{1}$ and $u_{2}$; the rate of Consistent Recommendation, which assesses if CRS recommends the same items given two user responses $u_{1}$ and $u_{2}$; and the rates of Diversity, which evaluate whether the recommended items, even if inconsistent, align with user preferences. We refer to it as Sensitivity when CRS provides inconsistent and inaccurate recommendations that do not align with user preferences.
3.1.2 Social Intelligence
It requires CRS to produce adequate social behavior for the recommendation during the conversations. As evidenced by the Media Equation Theory [^3] Users tend to engage with the machine in a manner that mirrors person-to-person conversations. Reeves and Nass (1996); Fogg (2003), users have high expectations for CRS to act cooperatively and be aware of user’s social needs during the conversation, facilitating the design of CRS with perceived humanness Jacquet et al. (2018, 2019). Our categorization encompasses two abilities.
  • Cooperation. CRS should follow the cooperative principle to achieve comfortable conversations in common social situations. This is accomplished by adhering to the four "Maxims of Conversation" Grice (1975, 1989), which form the basis for cooperative capability: 1) Manner. CRS should respond in a manner that is easily understood and clearly expressed. 2) Sincerity. CRS should communicate sincerely without deception, and ensure that its responses are backed by sufficient evidence. 3) Response Quality. CRS should provide the necessary level of information for the conversation without overwhelming the user with unnecessary details. 4) Relevance. CRS’s responses should contribute to identifying user preferences and making recommendations.
  • Social Awareness. CRS must meet users’ social expectations in practice, showing care, empathy, and establishing rapport with them Björkqvist et al. (2000). This facilitates the authenticity of CRS Neururer et al. (2018). To achieve this, a recent study Hayati et al. (2020) identified eight social strategies for building rapport with users, e.g., CRS could engage in self-disclosure and share its subjective opinion about a movie to establish a social connection with users.
We mainly resort to an LLM-based evaluator, except for sincerity evaluation (cf. Section 4.2). For sincerity evaluation, we aim to utilize objective metrics: the ratio of non-existent items and deceptive tactics. For non-existent items, we tally how many CRS-recommended items do not exist in the dataset. As for the deceptive tactics, we focus on the items accepted by the user and track how many of these do not align with user preferences. If the user accepts these items, we consider the CRS to be employing deceptive tactics, leading the user to believe that the items meet his preferences.
Abilities Descriptions Evaluation metrics
Quality
Provide precise recommendations
using minimal conversation turns
Computational metrics using the User Acceptance Rate
Reliability
Deliver robust and consistent recommendations
that account for contextual nuances
Computational metrics using the ratio of inconsistent recommendation
and the ratio of recommendation sensitivity
Cooperation
Follow cooperative principle
to achieve comfortable conversations
The average score of the Manner, Sincerity,
Response Quality, and Relevance
   1. Manner
Response should be easily understood
and clearly expressed
Ability-specific scoring
   2. Sincerity
Communicate sincerely,
without deception of pretense
Computational metrics using the ratio of deceptive tactics
and the ratio of non-existent items
   3. Response Quality
Provide the necessary level of information
without unnecessary details
Ability-specific scoring
   4. Relevance
Responses should contribute
to making recommendations
Ability-specific scoring
Social Awareness
Meet user social expectations,
establishing rapport with them
Ability-specific scoring
Identity
Self-aware of its identity
and operate within its designated scope
Computational metrics using Ratio of deceptive tactics
Coordination
Proficient in serving various and unknown users
without prior coordination
Computational metrics using the range and mean of
other ability-specific scores that are calculated among various users.
Table 1: Summary of the evaluation taxonomy, descriptions of abilities, and evaluation metrics in Concept. The LLM-based evaluator is used for ability-specific scoring, whereas computational metrics are used for the rest (cf. Table 18 in Appendix for details). We adjust the score to a scale of 1 to 5 when needed.
3.1.3 Personification
Personification requires CRS to perceive the identity of itself and the personality representation of users. This involves CRS being self-aware of its role as a conversational recommendation system tailored for the general public.
  • Identity. CRS should be self-aware of its identity and operate within its designated scope, differentiating itself from sales systems or other types. This ensures that CRS effectively carries out its identity by offering persuasive explanations to boost user acceptance Jannach et al. (2021); Zhou et al. (2022), avoiding resorting to misleading strategies (i.e., sales pitches with deceptive tactics). Otherwise, user trust and loyalty are hindered Gkika and Lekakos (2014). Using misleading strategies also violates the maxim of sincerity (cf. Section 3.1.2). We emphasize the need to integrate self-awareness into CRS to mitigate any deceptive behavior.
    We evaluate the identity by calculating the proportion of non-deceptive explanations. We also resort to the LLM-based evaluator to score the persuasiveness of recommendation explanations.
  • Coordination. CRS should be proficient in serving users possessing various personas without prior coordination. Attributing personality to a conversational AI ensures that its behaviors stand in agreement with the users’ expectations in a particular context Chaves and Gerosa (2021). This becomes particularly challenging for CRS, as it frequently serves users with diverse personas in real-world scenarios Thompson et al. (2004). As a result, CRS needs to exhibit various personalities and adapt its behavior to suit different users Katayama et al. (2019); Svikhnushina et al. (2021); Zhou et al. (2020). A crucial ability is to proficiently serve users without prior coordination.
    We create simulations of users with different personas and evaluate the performance of the CRS across all the abilities mentioned. To assess coordination, we initially calculate the range [^4] The range is more effective than the standard deviation in highlighting the variability of the CRS across different users. and mean of the CRS’s scores for each specific ability across different users. Then, we divide the range by the mean to determine the coordination score of the CRS for that specific ability. The overall coordination score is subsequently calculated as the average across all abilities.
4 Experiment and Evaluation
4.1 Experimental Setup
Dataset. We utilize GPT-3.5-16K-turbo to create personas and play the roles of the LLM-based user simulator and evaluator. Following Wang et al. (2023d), user preferences are defined using attributes from two benchmark datasets, i.e., Redial Li et al. (2018) and OpendialKG Moon et al. (2019). We report the statistical summary of the generated conversation dataset [^5] We will release all the conversation data and corresponding evaluation results when necessary.in Table 5.
Metrics Redial OpendialKG
KBRD BARCOR UNICRS CHATCRS KBRD BARCOR UNICRS CHATCRS
Recommendation Module Perspective Recall@1 0.02 0.22 0.13 0.41 0.12 0.03 0.15 0.37
Recall@10 0.23 1.37 1.09 2.27 0.98 0.94 1.28 3.23
Recall@25 0.57 3.23 2.44 4.95 1.94 2.07 3.06 8.20
Recall@50 1.13 5.69 4.58 8.85 3.53 3.43 5.81 15.14
SR@3 3.95 31.36 14.04 37.72 4.69 1.82 9.90 31.12
SR@5 4.39 35.53 15.68 40.90 14.19 3.52 17.45 37.24
SR@10 4.50 39.47 18.20 46.60 16.02 7.29 29.30 46.48
AT ($\downarrow$) 3.30 3.80 2.86 2.50 4.07 4.19 5.14 3.56
Conversation Module Perspective SR@3 20.18 27.52 35.20 52.63 6.51 17.71 14.58 26.30
SR@5 24.34 39.47 38.27 58.55 10.68 24.22 26.69 36.33
SR@10 29.39 50.66 43.42 62.39 12.37 35.16 45.31 44.40
AT ($\downarrow$) 2.07 2.87 3.02 3.23 3.97 5.88 5.00 3.74
User Perspective Acceptance Rate 0.33 1.43 0.33 70.83 0.39 0.65 0.26 64.32
AT ($\downarrow$) 8.01 5.62 7.67 4.75 5.33 6.40 5.00 4.69
Table 2: Recommendation quality evaluation (%) from three different perspectives. The average turn (AT) is calculated based on the corresponding conversation data with successful recommendations.
CRS Model. Following (Wang et al., 2023d), we present a comparative evaluation and analysis towards representative and SOTA CRS models, including KBRD Chen et al. (2019), BARCOR Wang et al. (2022a), UNICRS Wang et al. (2022b), and CHATCRS Wang et al. (2023d). Refer to Appendix A for implementation details.
Figure 2: An overview of the evaluation of off-the-shelf CRS models. The practical usability of those models are still hindered.
4.2 Evaluation Using Concept
Following the recent study on CRS Wang et al. (2023d), Concept resorts to an LLM-based user simulator and evaluator for cost-effective evaluation, together with computational metrics. We summary the taxonomy and evaluation metrics in Table 1. See Appendix A.2 for evaluation details.
Evaluation Process. The simulator creates simulations of users with diverse personas and preferences to interact with the CRS in producing conversation data. Afterwards, the evaluator assigns scores based on ability-specific scoring rubrics. Specifically, Concept considers the free-form chit-chat between the user simulator and CRS. To simulate real-world scenarios, the simulator has no access to its targeted items during the conversation. Any item that meets these preferences, such as having attributes completely consistent with or containing the simulator’s preferences, is considered a successful recommendation. During the conversation, Concept allows the simulator to describe their preferences in their own words, as in real-world situations, users may not use the exact terms defined in the pre-defined preference values. Finally, the conversation will end if the simulator accepts recommendations or if the conversation reaches the maximum number of turns. Finally, Concept utilizes both the LLM-based evaluator and computational metrics to assess the abilities of CRS. Note that the evaluator is utilized when corresponding computational metrics are not available.
LLM-based User Simulator. Our user simulator has unique personas and preferences. The personas are generated by prompting ChatGPT in a zero-shot manner, following (Wang et al., 2023a), while the preferences are defined using attributes from two benchmark datasets (refer to Section 4.1). In addition, Concept incorporates the Theory of Mind into our simulator to emulate human social cognition Fischer (2023). This is achieved by prompting the simulator to first assess its current mental state before generating responses, eliciting reflection on its predefined personality traits and social interactions. See Appendix A.2.2 for more implementation details on the simulator.
LLM-based Evaluator. Following previous studies Ye et al. (2023); Wang et al. (2023c), we employ the instance-wise evaluator. Building upon prior research Wang et al. (2023d); Liu et al. (2023d), we prompt the evaluator with fine-grained scoring rubrics to eliminate the scoring bias [^6] See Appendix D for our reliability analysis.. The evaluator assigns a score ranging from 1 to 5 to the conversation data using ability-specific rubrics, each accompanied by a corresponding description. For the generation of fine-grained rubrics, we follow the approach of previous works Saha et al. (2023a); Li et al. (2023b) and employ ChatGPT to produce a set of evaluation criterion, which are then refined by humans. See Appendix A.2.3 for details.
4.3 Evaluating off-the-shelf CRS models
Figure 2 presents an overview of the results across the six primary abilities, averaged across two benchmark datasets [^7] For results on each benchmark, refer to Appendix E.. These results suggest that CHATCRS has made significant progress in cooperation, social awareness, and recommendation quality while losing advantages in identity. We will continue the current discussion in subsequent sections and provide further analysis.
4.3.1 Recommendation-centric Evaluation
CHATCRS stands out as the leading CRS model, delivering higher-quality recommendations. As indicated in Table 2, most CRS models have demonstrated improved success rates (SR) from the conversation module’s perspective, indicating the effectiveness of optimizing conversation and recommendation modules simultaneously. However, we observed lower SR values for BARCOR on the Redial dataset compared to those from the recommendation module perspective, which is linked to the inclusion of non-existent items by its conversation module, such as "The Adventures of Milo and Ours" (should be ’Otis’) and "The Prestigige" (should be ’Prestige’). Despite this, their recall [^8] We only specifies user preferences, leading to a wide range of target items for users, resulting in a low recall value. and success rate (SR) values still fall short. In addition, the user acceptance rate of CRS models other than CHATCRS is notably low, possibly due to users’ intolerance towards poor recommendations and the absence of persuasive recommendation explanations, which we will investigate on in further studies. Instead, CHATCRS stands out as the leading CRS model, with significant performance improvements, conforming to recent findings Wang et al. (2023d). This may be attributed to two factors: 1) The text-embedding-ada-002 is much stronger in terms of translating conversation context and user preferences into embeddings, which helps enhance the recommendation module significantly. 2) CHATCRS are persuasive in convincing users to accept the recommendations. However, such a high user acceptance rate is largely based on deceptive tactics, a topic we will delve into later.
Figure 3: Reliability evaluation on CHATCRS, which is sensitive to contextual nuance.
Current CRS strives to offer reliable recommendations due to its sensitivity to contextual nuance. We gauge the model reliability by evaluating how off-the-shelf CRS models handle user responses that are semantically similar but expressed differently. Results of CHATCRS [^9] More results are in Table 12. are presented in Figure 3, which show promising action consistency of CHATCRS. It can make consistent recommendations when presented with two semantically similar user responses, achieving a consistency rate of over 99%. However, the average recommendation consistency rate is much lower, at only 51.58%, which means that 48.42% of the time, if the user alters their wording slightly, CHATCRS will recommend entirely different items. To further explore its impact, we conducted a thorough analysis of the relationship between the attributes of the recommended items and user preferences. We discovered that only about 12%-17% of the recommended items align with user preferences, despite being inconsistent. This may increase chances for relevant but less popular items. As a consequence, the majority of recommendation results do not align with user preferences, demonstrating the sensitivity of current CRS to contextual nuances and the negative impact on user experience.
4.3.2 Social-centric Evaluation
We break down cooperation ability into four maxims with an in-depth analysis, together with the social awareness evaluation. The results are depicted in Figure 4.
Figure 4: Evaluation on social-centric characteristics. CRS strives to self-express sincerely.
CHATCRS manifests politeness and conversational traits in terms of manner, quality, relevance, and social awareness. Different from other methods that lack a sense of topic or unable to maintain continuous conversations, CHATCRS has greatly improved its language abilities and social skills, which can be directly attributed to the impressive capabilities of ChatGPT in NLU, NLG, and empathy. However, there is still room for improvement in terms of social awareness, even for CHATCRS. For example, there are instances where CHATCRS recommends items generated from a few rounds ago. As evidenced by previous studies Portela and Granell-Canut (2017), keeping track of the conversational history is reported as an empathic behavior, leading to the rise of affection. This calls for the investigation on a more social-aware CRS.
Figure 5: Coordination evaluation on users with various personas (marked with dashed boxes). CHATCRS is hard to cater to diverse users. Others consistently show low performance levels that don’t vary much across different users.
Current CRS struggle to express genuine responses without hallucination or deceit. According to Fig. 4, sincerity scores of CRS models are far from satisfactory, which can be attributed to two factors: the recommendation hallucination and dishonest explanations. Regarding the hallucination, CRS tends to introduce non-existent items into conversations and promote them to users. Even in the case of CHATCRS, its responses still contain 5.18% of non-existent items on Redial and 7.42% on OpendialKG. As for dishonest explanations, CRS tends to utilize persuasive language to mislead users into accepting recommendations by providing false explanations in movie plots and attributes. This problem is even more severe in CHATCRS, where, on average, approximately 62.09% of the explanations do not meet the requirements of sincerity (cf. Section 4.3.3, Fig. 6).
Dataset KBRD BARCOR UNICRS CHATCRS
Redial 1.02 1.55 1.08 4.66
OpendialKG 1.00 1.25 1.06 4.48
Avg. 1.01 1.40 1.07 4.57
Table 3: Results of persuasiveness scores.
4.3.3 Personification-centric Evaluation
Lacking self-awareness, CRS models often offers persuasive yet dishonest explanations. As in Table 3, most CRS models aim for being persuasive in their recommendations. Explanations may sound lackluster, insipid, or feeble, simply stating that "this is a good movie," or making absurd claims such as "Jumanji (2017) is about a man who is a human!". These recommendations become less acceptable for users. In contrast, CHATCRS consistently furnishes explanations for its recommendations, employing text-based logical reasoning to enhance comprehensibility. This accounts for its high user acceptance rate. However, as in Figure 6, although highly convincing, the explanations from CHATCRS often include illusory details, leading users to mistakenly believe that these items align with their preferences. Taking OpenredialKG for example, 75.10% of accepted items do not align with user preferences. Given that the majority of CRS models employ reinforcement learning, or RLHF, this could cause problems such as reward hacking and misspecification Pan et al. (2022). Such problems can drive CRS to acquire deceitful behavior. Considering the above disadvantages, the urgency is to create an identity-aware CRS that can deliver both persuasive and honest explanations.
Figure 6: Identity and sincerity evaluation. CHATCRS offers persuasive yet dishonest explanations.
Lacking proficiency in prior coordination, CRS models often fails to cater different user needs. As in Figure 5, most CRS models, except for CHATCRS, show poor performance in sensing the variation of users. In contrast, CHATCRS outperforms others in general and is more sensitive to different users, providing high-quality recommendations. In particular, CHATCRS can properly deal with users’ negative emotions, such as bored, confused, or disappointed. Interestingly, these people tend to show low acceptance of recommendations, i.e. they possess a higher acceptance threshold for recommendation quality. Such observations drives us in developing a CRS that can dynamically adjust its recommendation strategy to align with the unique personas of each user. Apart from that, we also examine the interaction behavior of CHARCRS with different users. For instance, CHATCRS adopts sales pitches with deceptive tactics to persuade optimistic users to accept recommendations. However, for pessimistic users, CHATCRS tends to provide persuasive and honest recommendation explanations. This somewhat reveals a bias in CHATCRS’s recommendation strategy towards different user groups, which is a flaw that needs to be rectified in future work.
Reliability of Concept. We have established fixed values for the temperature and seed parameters to guarantee replicability of our LLM-based simulator and evaluator. We show that Concept scoring is unbiased and aligns with the results of human evaluations (cf. Appendix D).
5 Conclusion
We regard CRS as a social issue, instead of just a technical problem, considering its major goal of persuading users to accept the recommendations. Hence, we revisit current evaluation protocols and pioneer works to provide a comprehensive survey on understanding the impact of CRS’s characteristics in terms of both pros and cons. We then come up with a new evaluation protocol– Concept, which considers both system- and user-centric factors for addressing personalized user needs in particular, with the hope of enhancing user experience in CRS. Such factors include three key characteristics, which are further divided into six primary abilities. In addition, we suggest that even enhanced by "omnipotent" LLMs, CRS lacks practical usability. We pinpoint several significant limitations of current CRS models, including e.g. the incapacity to express itself sincerely, a lack of identity-awareness, low reliability towards contextual nuance, and poor coordination for diverse users. Through Concept, we aim for providing CRS developers and practitioners with a guidance to draw meaning insights from our inclusive and fine-grained evaluation directions.
6 Limitations
The use of LLM-based user simulators and evaluators is a double-edged sword. It is a labor-intensive and effective approach, but may suffer from weak robustness by its nature. Although we have adopted some strategies to improve the robustness as much as possible following previous works, the robustness may still be hampered due to the potential uncertainty of prompt engineering. Additionally, another aspect that could improve the Concept is to generate more conversation data, run the LLM-based evaluation many times with different seeds, and report more statistically significant results. However, budgeting is always a factor to consider. In this case, it is important to propose a user simulator and evaluator based on an open-source small model that has similar capabilities to ChatGPT. Finally, our current work does not evaluate the attribute-based CRS Lei et al. (2020), as this type of research often ignores the ability to engage in smooth conversations and instead focuses solely on accurately providing recommendations to users within minimal conversation turns. In this case, evaluating attribute-based CRS models seems unfair. We highlight the importance of combining attribute-based and dialog-based CRS studies to create a more holistic CRS, taking into account its practical usability.
Appendix A Implementation Details
We conduct all our experiments using a single Nvidia RTX A6000, and we implement our codes in PyTorch. The core of our code framework is built upon open-source code from the latest research on CRS [^10] https://github.com/RUCAIBox/iEvaLM-CRS Wang et al. (2023d). In order to guarantee replicability, we have established fixed values for the Temperature and Seed parameters of ChatGPT (i.e., GPT-3.5-16K-turbo), setting the Temperature to 0 and the Seed to 42.
A.1 Implementation of CRS Models
We evaluated all open-source CRS models in its code using their checkpoints, including the KBRD Chen et al. (2019), BARCOR Wang et al. (2022a), UNICRS Wang et al. (2022b), and CHATGPT-based model Wang et al. (2023d) which represents the current SOTA, incorporating text-embedding-ada-002 Neelakantan et al. (2022) for modeling the semantic embeddings.
  • KBRD Chen et al. (2019) bridges the recommendation module and the Transformer-based conversation module via knowledge propagation.
  • BARCOR Wang et al. (2022a) is a unified framework based on BART Lewis et al. (2020), which implements the recommendation and response generation tasks in a single model.
  • UNICRS Wang et al. (2022b) is a unified framework based on DialoGPT Zhang et al. (2020), with a semantic fusion module to enhance the semantic association between conversation history and knowledge graphs.
  • CHATCRS Wang et al. (2023d) is the SOTA CRS model, which incorporates ChatGPT for the conversation module and the text-embedding-ada-002 Neelakantan et al. (2022) to enhance the recommendation module.
See the prompts in the original paper for the implementation of CHATGPT-based model. Note that we feed the top 5 items from the recommendation module into the prompts of ChatGPT for re-ranking and generating responses. Also, the UNICRS utilizes DialoGPT-small as the backbone, while the BARCOR utilizes BART-base with a 2-layer encoder and decoder, following Wang et al. (2023d).
Statistics Num
# Conversations 6720
Max Turns 10
Avg. Turns 8.92
Persona Types 12
Table 5: Statistics of the generated conversation dataset.
A.2 Implementation of Concept
Concept resorts to an LLM-based user simulator and evaluator for cost-effective evaluation, together with fine-grained ability-specific scoring rubrics. Table 4 reveals the statistics of our generated conversation dataset.
A.2.1 Evaluation Process
We used user simulators to interact and converse with different CRSs to produce a conversation dataset. In the experiments, the simulator simulated 12 different personas and 4 different age groups using ChatGPT. Specifically, Concept considers the free-form chit-chat between the user simulator and CRS. To simulate real-world scenarios, the simulator has no access to its targeted items during the conversation. Any item that meets these preferences, such as having attributes completely consistent with or containing the simulator’s preferences, is considered a successful recommendation. During the conversation, Concept allows the simulator to describe their preferences in their own words, as in real-world situations, users may not use the exact terms defined in the pre-defined preference values. During the conversation between CRS and the user, we recorded the recommendation results of each round of the recommendation system and the results recommended by the conversational agent, in order to evaluate the success of the recommendations from different perspectives. Finally, the conversation will end if the simulator accepts recommendations or if the conversation reaches the maximum number of turns. If the user chooses to accept the recommendation from the conversational agent, the user is required to add ’[END]’ at the end of his response to indicate the end of the conversation. These diverse users generated a total of 6720 conversations, which were used to evaluate the performance of different CRS in more realistic scenarios.
Afterward, Concept utilizes both the LLM-based evaluator and computational metrics to assess the abilities of CRS, as outlined in Table 1. Note that the evaluator is utilized when corresponding computational metrics are not available. Specifically, we employ the evaluator to assess the cooperation ability, social awareness, and persuasiveness of recommendation explanations of CRS. This involves prompting the evaluator with fine-grained, ability-specific scoring rubrics Saha et al. (2023b), which are generated by the LLM and then refined by humans. For more details, refer to Appendix A.2.3 and Appendix A.2.4.
A.2.2 LLM-based User Simulator
Our user simulator has unique personas and preferences. Specifically, the personas are generated by prompting ChatGPT in a zero-shot manner, following (Wang et al., 2023a), while the preferences are defined using attributes from two benchmark datasets, Redial and OpendialKG.
Personas. We generate a persona list as the starting point by prompting ChatGPT in a zero-shot manner. Inspired by Wang et al. (2023a), we first instruct the ChatGPT to return 20 distinct personas and their corresponding descriptions, such as curiosity, and then filter out duplicates. We utilize the generated persona list, so it is highly likely that ChatGPT is familiar with these personas and would contain knowledge and information about them. As a result, we obtain 12 unique personas, simulating diverse sentiments may encountered when using a conversational recommendation system (namely, Anticipation, Boredom, Fusion, Curiosity, Delight, Disappointment, Exceptions, Frustration, Independence, Surprise, Trust, Satisfaction), and each type of user may belong to different age groups (namely, Adults, Children, Senior, and Teens). We assigned different attribute groups as movie preferences for each user type. These users with varying emotions, age groups, and movie preferences then interacted with the CRS. If the attribute group of the movie recommended by the CRS is equal to or includes the user’s target attribute group, it is considered a correct recommendation. For the Redial dataset, each user type corresponds to 76 conversation data with one CRS, resulting in a total of 912 conversation data across all user types and one CRS. Similarly, for the OpendialKG dataset, each user type generates 64 conversation data with each CRS, resulting in a total of 768 conversation data across all user types and one CRS. Since we tested 4 different CRSs, we collected a total of 6720 conversation data. Refer to Table 17 for details.
Raw Attribute ChatGPT-adjusted Attributes
Redial
action thrilling and adrenaline-pumping action movie
adventure exciting and daring adventure movie
animation playful and imaginative animation
biography inspiring and informative biography
comedy humorous and entertaining flick
crime suspenseful and intense criminal film
documentary informative and educational documentary
drama emotional and thought-provoking drama
family heartwarming and wholesome family movie
fantasy magical and enchanting fantasy movie
film-noir dark and moody film-noir
game-show entertaining and interactive game-show
history informative and enlightening history movie
horror chilling, terrifying and suspenseful horror movie
music melodious and entertaining music
musical theatrical and entertaining musical
mystery intriguing and suspenseful mystery
news informative and current news
reality-tv dramatic entertainment and reality-tv
romance romantic and heartwarming romance movie with love story
sci-fi futuristic and imaginative sci-fi with futuristic adventure
short concise and impactful film with short story
sport inspiring and motivational sport movie
talk-show informative and entertaining talk-show such as conversational program
thriller suspenseful and thrilling thriller with gripping suspense
war intense and emotional war movie and wartime drama
western rugged and adventurous western movie and frontier tale
OpendialKG
Action adrenaline-pumping action
Adventure thrilling adventure
Sci-Fi futuristic sci-fi
Comedy lighthearted comedy
Romance heartwarming romance
Romance Film emotional romance film
Romantic comedy charming romantic comedy
Fantasy enchanting fantasy
Fiction imaginative fiction
Science Fiction mind-bending science fiction
Speculative fiction thought-provoking speculative fiction
Drama intense drama
Thriller suspenseful thriller
Animation colorful animation
Family heartwarming family
Crime gripping crime
Crime Fiction intriguing crime fiction
Historical drama captivating historical drama
Comedy-drama humorous comedy-drama
Horror chilling horror
Mystery intriguing mystery
Table 6: Illustration on ChatGPT-adjusted attributes. We provide the user simulators with adjusted attributes to prevent them from revealing their target attributes. In real-world situations, users may not always use the same words as those used during model training to express their preferences.
Preferences. The current evaluation method for CRS is not reflective of real-world situations, as it assumes that every user knows the target item Wang et al. (2023d). To tackle this issue, our user simulators only have access to their preferences. The user would only choose to accept the recommendation when the CRS provides explanations for the recommendations and makes the users feel that the recommended items match their preferences. To achieve this, the preferences are defined using attributes from two benchmark datasets, Redial Redial Li et al. (2018) and OpendialKG Moon et al. (2019). For each dataset, every movie has a feature group made up of one or more attributes. In the case of the Redial dataset, we conducted experiments using feature groups containing 3 attributes and excluded less common groups. Ultimately, we retained the 19 most prevalent attribute groups for the study, with each group corresponding to at least 50 different movies. For the OpendialKG dataset, the issue of less common attributes is more pronounced. Initially, we selected the most prevalent attributes (each corresponding to at least 100 movies) and then kept the 16 most common attribute groups for experimentation.
Simulation via Prompts. We prompt the simulator with different personas using the persona descriptions generated by ChatGPT (cf. Table 7.). Concept incorporates the Theory of Mind into our simulator to emulate human social cognition Fischer (2023). This is achieved by prompting the simulator to first assess its current mental state before generating responses, encouraging reflection on its predefined personality traits and social interactions. Refer to Appendix F for prompts. Additionally, there is a maximum of 10 turns allowed in the conversation, and it will only conclude when the simulated user accepts the recommendation. To prevent the simulated users from directly stating the same attribute preferences as the pre-defined values (i.e., the attribute group) when asked about their preferences by the CRS, we include adjectives before each attribute during the simulation, and we allow the user to describe his/her preference using their own words. This approach is reasonable because in real scenarios, users may not use the exact words as the pre-defined attribute values during the conversation. We have summarized the adjusted values for each attribute in Tables 6, achieved by prompting ChatGPT.
A.2.3 LLM-based Evaluator
We summarize the evaluation method for each characteristic / ability in Table 1. We utilize an LLM-based evaluator to evaluate characteristics or abilities when corresponding computational metrics are not available. These include the abilities of the Manner, Response Quality, and Relevance in Cooperation, Social Awareness, and ability of persuasiveness in Identity.
Based on the findings of previous studies Ye et al. (2023); Wang et al. (2023c, d), we utilize an instance-wise evaluator to conduct detailed assessments. Expanding on earlier research Ye et al. (2023); Wang et al. (2023d); Liu et al. (2023d), we task the evaluator with using fine-grained scoring rubrics to mitigate scoring bias (see Section D for our reliability analysis). The evaluator assigns scores ranging from 1 to 5 based on ability-specific score rubrics, each accompanied by a corresponding description, using the conversation data. For the generation of fine-grained scoring rubrics, we follow the approach of previous works and employ ChatGPT to produce a set of evaluation criteria Saha et al. (2023a). This set of criteria serves as a starting point for human refinement Li et al. (2023b). Prior to assigning a score, we require the evaluator to provide a rationale, drawing inspiration from the effectiveness of CoT prompting Wei et al. (2022); Ye et al. (2023). For implementation details on the evaluator, refer to Appendix F.
Persona Templates (The Input of ChatGPT Paraphraser) ChatGPT-paraphrased Persona Descriptions
Emotion=Boredom
Age group=Adults
you are a person that are easy to be Boredom.
This means that your are Feeling uninterested
or uninspired by the recommended movie choices.
Also, you are a Adults person
You are easily bored,
feeling uninterested or uninspired by the recommended movie choices.
As an adult, you seek movies that can captivate your attention.
Emotion=Anticipation
Age group=Children
you are a person that are easy to be Anticipation.
This means that your are Looking forward to
watching recommended movies and experiencing new stories.
Also, you are a Children person
You are filled with anticipation,
looking forward to watching recommended movies and experiencing new stories.
As a child, you enjoy the excitement of discovering new films.
Table 7: Persona description generation. We start by converting each combination of personality and age into a single sentence using a template, and then we use ChatGPT to rephrase it into a natural-sounding sentence. Refer to Appendix F for Prompts.
A.2.4 Computational Metrics
We introduce the computational metrics for evaluating the remaining characteristics or abilities. These include the abilities of quality and reliability in Recommendation intelligence, the ability of sincerity (Cooperation) in Social intelligence, and the abilities of identity and coordination in Personification. We conclude them as follows.
  • Quality. We consider computational metrics for automatic evaluation following previous works Wang et al. (2023d); Zhang et al. (2023); Yu et al. (2023), including $Recall@k(k=1,10,25,50)$, recommendation success rate $(SR@k,k=3,5,10)$, user acceptance rate ($AR$), and average turns ($AT$) required to reach the successful recommendation. Note that we determine user acceptance of the recommendation by checking for the presence of the special token ’[END]’ in the user’s response.
  • Reliability. We generate sets of user response pairs with similar meanings using ChatGPT paraphrasing. This enables us to evaluate how reliably the CRS performs with contextual nuances. Formally, given the same conversation history and a user response pair $u_{1}$ and $u_{2}$, we define four metrics to measure reliability: the rate of Consistent Action, indicating whether the CRS constantly provides recommendations based on $u_{1}$ and $u_{2}$; the rate of Consistent Recommendation, which assesses if the CRS recommends the same items given two user responses $u_{1}$ and $u_{2}$; and the rates of Diversity, which evaluate whether the recommended items, even if inconsistent, align with user preferences. We refer to it as Sensitivity when the system provides inconsistent and inaccurate recommendations that do not align with user preferences.
  • Sincerity. We value the ratio of non-existent items and deceptive tactics. For non-existent items, we have tallied how many items recommended by the CRS do not exist in the dataset. As for the deceptive tactics, we focus on the items accepted by users and tally how many of these do not align with the users’ pre-defined preferences. If users accept these misleading items, we consider the CRS to be employing a deceptive tactics, leading users to believe that the recommended items meet their preferences.
  • Identity. CRS should be self-aware of its identity and operate within its designated scope, differentiating itself from sales systems. We evaluate this by assessing the sincerity of explanations, i.e., the deceptive tactics.
  • Coordination. We simulate users with different personas and assess how the CRS performs across all previously mentioned abilities. To qualify the coordination score, we initially computed the Range and mean of the CRS model’s scores across various users, based on the different abilities mentioned earlier. It’s worth noting that the Range is more effective than the standard deviation in highlighting the variability of the CRS model across different users. Subsequently, we divided the Range by the mean to derive the Coordination score of the CRS for that specific ability. The overall Coordination score of the CRS is then calculated as the average across all abilities.
CRS Models System-centric User-centric Quantitative Implementations
Recommendation intelligence Social intelligence Personification
Quality Reliability Cooperation
Social
Awareness
Identity Coordination
Realistic
Environment
Well-established
CRS
CRS-Que Jin et al. (2023)
CRS-UX Jin et al. (2021)
US Siro et al. (2023)
INSPIRED Hayati et al. (2020)
iEval Wang et al. (2023d)
Concept (ours)
Table 8: Differences between our Concept and existing evaluation protocols. Concept focuses on inclusively assessing the foundational abilities that impact the practical use of CRS, bringing together system- and user-centric factors. The checkmark with cross means that only certain evaluation aspects are being addressed.
Appendix B More Related Works
Conversational Recommender System (CRS). CRS could be regard as a recommendation system that can elicit the dynamic preferences of users and take actions based on their current needs through real-time multi-turn conversations. Together with the conversation intelligence and recommendation techniques, CRS have evolved from being mere tools to autonomous agents with their own goals and strategies for interacting with users, especially in the era of LLM. In the early stages, CRS was constrained by predefined interaction options Tou et al. (1982); Smyth and McGinty (2003). However, the introduction of an advanced conversation module expanded the interaction capabilities of CRS to natural language Chen et al. (2019); Wang et al. (2022a), enabling users to communicate more flexibly and effectively with the system. Recent studies also witnessed the emergence of advanced capabilities in CRS, including proactivity Deng et al. (2023), conversation context modeling Li et al. (2023c); Wang et al. (2023d); Liu et al. (2023c), preference modeling Qian et al. (2023), and social awareness Hayati et al. (2020). While significant progress has been achieved, the usefulness of current CRS in practice remains to be questioned Jannach and Manzoor (2020); Jannach et al. (2021). This question is closely tied to how to evaluate CRS. However, there still lacks a comprehensive evaluation and understanding of how CRS performs in practice.
CRS Evaluation. The practical effectiveness of CRS has always been questioned Jannach and Manzoor (2020); Jannach et al. (2021). Addressing this concern necessitates an inclusive evaluation protocol to assess the strengths, weaknesses, and potential risks of CRS. The success of CRS in practice relies primarily on social characteristics, not just technical ones. However, most evaluation protocols primarily focus on superficial and limited evaluation aspects (in isolation) from system-centric perspectives, such as lexical diversity and perplexity of responses Ghazvininejad et al. (2018); Chen et al. (2019), conversational fluency, relevance, informativeness Wang et al. (2022b, a), recommendation effectiveness and efficiency Wang et al. (2023d); Jin et al. (2019); Wärnestål (2005), and rapport building Hayati et al. (2020). Since they fail to reveal the inadequacy of CRS from user-centric perspectives, these fragmented evaluation protocols are underdeveloped to constitute a standard benchmark and conduct an inclusive evaluation. Recent efforts have attempted to systematically summarize and evaluate the characteristics of CRS, but they have mainly centered on establishing evaluation protocols through person-to-person conversation analysis and questionnaire interviews, lacking quantitative observations on existing CRS models Jin et al. (2023, 2021); Jannach (2022); Siro et al. (2023) and failing to systematically evaluate system-centric characteristics. Refer to Table 8 for better understanding our differences. In summary, existing protocols are considered too superficial and inadequate to uncover the shortcomings of CRS Jannach et al. (2021); Jannach (2023). To this end, our Concept brings together system- and user-centric perspectives, enabling a more profound and inclusive evaluation protocol and ultimately contributing to the development of more practical CRS Li et al. (2023a).
LLM as User Simulator. Manual evaluation often requires the user to constantly interact with CRS, which are labor-expensive and generally only feasible in an industry lab Huang et al. (2023); Gao et al. (2021). To this end, LLM-based user simulators continue to be a rapidly evolving research area due to their labor-intensive nature and effectiveness Kocmi and Federmann (2023); Terragni et al. (2023); Tang et al. (2023); Cheng et al. (2023); Afzali et al. (2023). A recent study on CRS also features interactive evaluation with LLM-based user simulation Wang et al. (2023d), experimentally demonstrating its effectiveness and reliability as a reliable alternative to humans. Taking inspiration from these studies, Concept also harnesses an LLM-based user simulator to interact with the CRS. Moreover, we equip our simulator with elaborately designed instructions and the Theory of Mind to mimic human social cognition Fischer (2023). This encourages the simulator to reflect on its own predefined personality traits.
LLM as Evaluator. A LLM-based evaluator could be seen as a combination of a LLM and its prompting strategy Zeng et al. (2023); Cohen et al. (2023); Chan et al. (2023); Wang et al. (2023b); Liu et al. (2023b). Considering its flexibility, the instance-wise evaluator is widely adopted Ye et al. (2023); Wang et al. (2023c); Liu et al. (2023a). Importantly, previous findings highlight that offering detailed scoring rubrics or criteria contributes to achieving consistent and aligned evaluations with human assessments Liu et al. (2023d). These findings are further supported by recent work on CRS Wang et al. (2023d), resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. Inspired by them, we involve ability-specific scoring rubrics to gain reliable and aligned evaluations. We also encourage the evaluator to generate a rationale before assigning a score, further ensuring reliability Ye et al. (2023).
Appendix C Characteristics Categorization
We consolidate both system- and user-centric factors as the starting point. Interdisciplinary research on conversational AI Chaves and Gerosa (2021); Reeves and Nass (1996); Fogg (2003) groups its characteristics that impact the user experience into three types. Concept then tailors them for CRS, which conceptualizes expected characteristics into six primary abilities for the inclusive CRS evaluation. In Table 8, we elaborate the inclusive CRS evaluation protocol, with the special focus on the definitions of each abilities and skills, and how we evaluate them.
Attribute Group BARCOR CHATCRS KBRD UNICRS
Redial
[’action’, ’adventure’, ’animation’] 1.77 4.33 1.13 1.38
[’action’, ’adventure’, ’comedy’] 1.94 4.31 1.21 1.33
[’action’, ’adventure’, ’drama’] 1.63 4.19 1.08 1.40
[’action’, ’adventure’, ’fantasy’] 1.85 4.31 1.19 1.46
[’action’, ’adventure’, ’sci-fi’] 1.94 4.27 1.10 1.29
[’action’, ’adventure’, ’thriller’] 1.83 4.44 1.15 1.27
[’action’, ’crime’, ’drama’] 1.79 4.29 1.15 1.31
[’action’, ’crime’, ’thriller’] 1.83 4.42 1.19 1.40
[’adventure’, ’animation’, ’comedy’] 1.92 4.25 1.10 1.27
[’adventure’, ’comedy’, ’family’] 1.92 4.06 1.25 1.52
[’biography’, ’crime’, ’drama’] 1.53 4.21 1.08 1.23
[’biography’, ’drama’, ’history’] 1.65 4.63 1.10 1.27
[’comedy’, ’drama’, ’family’] 1.92 4.29 1.04 1.50
[’comedy’, ’drama’, ’romance’] 1.77 4.31 1.19 1.44
[’crime’, ’drama’, ’thriller’] 1.81 4.42 1.10 1.40
[’drama’, ’horror’, ’mystery’] 1.83 4.38 1.08 1.40
[’horror’, ’mystery’, ’thriller’] 2.08 4.17 1.10 1.40
[’crime’, ’drama’, ’mystery’] 1.83 4.46 1.10 1.29
[’action’, ’comedy’, ’crime’] 1.56 4.13 1.06 1.29
Avg.+-Std. 1.81+0.14 4.31+0.13 1.13+0.05 1.36+0.08
OpendialKG
[’action’, ’adventure’, ’fantasy’] 1.58 4.19 1.02 1.15
[’action’, ’adventure’, ’sci-fi’] 1.46 4.40 1.08 1.13
[’action’, ’adventure’, ’thriller’] 1.67 4.29 1.00 1.23
[’comedy’, ’drama’, ’romance’] 1.73 3.96 1.04 1.10
[’crime’, ’drama’, ’thriller’] 1.56 4.63 1.00 1.15
[’horror’, ’mystery’, ’thriller’] 1.58 4.33 1.02 1.19
[’Adventure’, ’Animation’,
’Comedy’, ’Family’]
1.56 4.31 1.04 1.17
[’Comedy’, ’Romance’, ’Romance Film’] 1.46 3.94 1.02 1.15
[’Action’, ’Adventure’, ’Sci-Fi’, ’Thriller’] 1.46 4.40 1.00 1.08
[’Fantasy’, ’Fiction’, ’Science Fiction’,
’Speculative fiction’]
1.65 3.48 1.06 1.08
[’Drama’, ’Historical period drama’,
’Romance’, ’Romance Film’]
1.58 3.88 1.00 1.15
[’Comedy’, ’Comedy-drama’, ’Drama’] 1.54 3.85 1.00 1.13
[’Action’, ’Crime’, ’Drama’, ’Thriller’] 1.65 3.50 1.04 1.19
[’Action’, ’Adventure’, ’Fantasy’, ’Sci-Fi’] 1.50 3.77 1.04 1.15
[’Crime’, ’Crime Fiction’, ’Drama’, ’Thriller’] 1.50 4.27 1.02 1.15
[’Comedy’, ’Romance’,
’Romance Film’, ’Romantic comedy’]
1.65 3.71 1.00 1.13
Avg.+-Std. 1.57+0.08 4.06+0.33 1.02+0.02 1.14+0.04
Table 9: Overall performance evaluation when recommending items with various attributes
C.1 Recommendation Intelligence
It requires CRS to learn from conversations and evolve toward making recommendations as the conversation advances. Taxonomy in Chaves and Gerosa (2021) highlights conversational intelligence, which pertains to the AI’s ability to facilitate conversations [^11] We evaluate CRS’s conversational proficiency within the context of social intelligence.. We highlight CRS’s capacity to model user preferences from conversations and create recommendations. As a basic function from system-centric perspectives, CRS needs to produce high-quality and reliable recommendations to fit user preferences based on the conversation history Chen et al. (2019); Ma et al. (2020); Zhou et al. (2021). In this regard, our categorization encompasses two primary abilities.
Quality. As the basic requirements, CRS should provide precise recommendations using minimal conversation turns, which are the crucial aspects that influence user satisfaction Siro et al. (2023); Gao et al. (2021). Apart from evaluating the quality of items recommended by the recommendation module and the conversation module [^12] Conversation module can conduct recommendation Chen et al. (2019), motivated us to gauge its quality. separately, we emphasize the importance of user acceptance rate, reflecting the practical effectiveness and relevance of the recommendations in practice.
Specially, we consider computational metrics for automatic evaluation following previous works Wang et al. (2023d); Zhang et al. (2023); Yu et al. (2023), including $Recall@k(k=1,10,25,50)$, recommendation success rate $(SR@k,k=3,5,10)$, user acceptance rate ($AR$), and average turns ($AT$) required to reach the successful recommendation. Note that we determine user acceptance of the recommendation by checking for the presence of the special token ’[END]’ in the user’s response.
Reliability. CRS should deliver robust and consistent recommendations that account for contextual nuances. In practical situations, users with similar preferences may express themselves differently. It would be frustrating to recommend divergent items for two semantically similar inputs Oh et al. (2022). This not only diminishes the user experience but can also be disruptive or even perilous for critical applications in healthcare and finance Tran et al. (2021); Zibriczky12 (2016). In less critical scenarios, if two recommended items are inconsistent but align with user preferences, they are often viewed as diverse recommendations, creating more opportunities for relevant but less popular items Yang et al. (2021). We also consider this when evaluating.
We generate sets of user response pairs with similar meanings using ChatGPT paraphrasing. This enables us to evaluate how reliably the CRS performs with contextual nuances. Formally, given the same conversation history and a user response pair $u_{1}$ and $u_{2}$, we define four metrics to measure reliability: the rate of Consistent Action, indicating whether the CRS constantly provides recommendations based on $u_{1}$ and $u_{2}$; the rate of Consistent Recommendation, which assesses if the CRS recommends the same items given two user responses $u_{1}$ and $u_{2}$; and the rates of Diversity, which evaluate whether the recommended items, even if inconsistent, align with user preferences. We refer to it as Sensitivity when the system provides inconsistent and inaccurate recommendations that do not align with user preferences.
Age Group BARCOR CHATCRS KBRD UNICRS
OpendialKG
Adults 1.54 3.94 1.03 1.15
Children 1.58 4.14 1.03 1.16
Seniors 1.54 3.99 1.01 1.12
Teens 1.62 4.15 1.03 1.15
Redial
Adults 1.84 4.38 1.11 1.32
Children 1.76 4.33 1.15 1.39
Seniors 1.79 4.23 1.10 1.38
Teens 1.86 4.29 1.14 1.36
Table 10: Overall performance evaluation when dealing with users of various ages.
C.2 Social Intelligence
It requires CRS to produce adequate social behavior for the recommendation during the conversations. Taxonomy in Chaves and Gerosa (2021) highlights social intelligence, which focuses on habitual social protocols. As evidenced by the Media Equation Theory Reeves and Nass (1996); Fogg (2003), users tend to engage with the machine in a manner that mirrors person-to-person conversations. They have high expectations for the machine to adhere to the cooperative principle Nass et al. (1996), and instinctively expect the machine to demonstrate empathy Björkqvist et al. (2000); Hayati et al. (2020) in order to facilitate comfortable conversations in social situations. These highlight the need for CRS to act cooperatively and be aware of the user’s social needs during the conversation, facilitating the design of CRS with perceived humanness Neururer et al. (2018); Jacquet et al. (2018, 2019). In this regard, our categorization encompasses two primary abilities, including cooperation and social awareness.
Cooperation. CRS should follow the cooperative principle to achieve effective and comfortable conversations in common social situations Reeves and Nass (1996); Fogg (2003); Chaves and Gerosa (2021). The cooperative principle guides individuals to communicate in a way that is helpful and contributes to the overall success of the conversation Nass et al. (1996). This is achieved by following four “Maxims [^13] Maxim is a phrase or saying that includes a rule or moral principle about how one should behave. of Conversation" Grice (1975, 1989); SARI (2020), which serve as a foundation for refining the cooperative ability of CRS: 1) Manner. CRS should respond in a manner that is easily understood and clearly expressed. 2) Sincerity. CRS should communicate sincerely, without deception or pretense, and ensure that its responses are backed by sufficient evidence. 3) Response Quality. CRS should provide the necessary level of information for the conversation without overwhelming the user with unnecessary details. 4) Relevance. CRS should ensure that its responses contribute to identifying user preferences and making recommendations.
Social Awareness. CRS needs to meet users’ social expectations in practice, showing care, empathy, and establishing rapport with them Björkqvist et al. (2000). This facilitates the authenticity of CRS Neururer et al. (2018). To achieve this, a recent study Hayati et al. (2020) identified eight social strategies for building rapport with users. For instance, CRS could engage in self-disclosure and share its subjective opinion about a movie to establish a social connection with users.
As for the evaluation implementations, We mainly resort to an LLM-based evaluator with elaborately designed instructions and ability-specific score rubrics (ranging from $1-5$) for evaluation (cf. Appendix A.2.3). As for the sincerity evaluation, we value the ratio of non-existent items and deceptive tactics. For non-existent items, we have tallied how many items recommended by the CRS do not exist in the dataset. As for the deceptive tactics, we focus on the items accepted by users and tally how many of these do not align with the users’ pre-defined preferences. If users accept these misleading items, we consider the CRS to be employing a deceptive tactics, leading users to believe that the recommended items meet their preferences.
Personas BARCOR CHATCRS KBRD UNICRS
Redial
Anticipation 1.76 4.91 1.24 1.39
Boredom 1.72 3.16 1.05 1.38
Confusion 1.84 3.49 1.13 1.32
Curiosity 1.86 4.82 1.16 1.41
Delight 1.78 4.47 1.14 1.38
Disappointment 1.82 3.33 1.08 1.33
Excitement 1.93 4.96 1.14 1.39
Frustration 1.68 4.67 1.07 1.26
Indifference 1.78 3.92 1.07 1.26
Satisfaction 1.83 4.46 1.17 1.49
Surprise 1.88 4.89 1.14 1.41
Trust 1.86 4.62 1.13 1.29
OpendialKG
Anticipation 1.69 4.67 1.05 1.16
Boredom 1.58 2.94 1.06 1.14
Confusion 1.38 3.38 1.00 1.11
Curiosity 1.63 4.58 1.05 1.08
Delight 1.58 4.00 1.02 1.17
Disappointment 1.56 3.00 1.02 1.14
Excitement 1.52 4.59 1.00 1.28
Frustration 1.47 4.38 1.00 1.05
Indifference 1.56 4.08 1.03 1.09
Satisfaction 1.63 4.13 1.03 1.19
Surprise 1.69 4.48 1.05 1.16
Trust 1.58 4.45 1.00 1.16
Table 11: Overall performance of CRS when dealing with users of various personas
C.3 Personification
It requires CRS to perceive the identity of itself and the personality representation of users. Taxonomy in Chaves and Gerosa (2021) highlights personification of conversational AI in terms of identity and personality, which value the different traits of conversational AI. This requires CRS to have self-awareness of its role as a conversational recommendation system, distinguishing itself from sales systems or other types of systems. This helps CRS operate within its intended scope and purpose, ensuring that it effectively fulfills its role by providing persuasive and honest explanations to encourage acceptance Jannach et al. (2021); Zhou et al. (2022) for diverse users Thompson et al. (2004). In this regard, our categorization encompasses two primary abilities.
Identity. CRS should provide persuasive and honest explanations to boost the acceptance of the recommendations. In real-world situations, users often have little knowledge about the recommended items. A persuasive explanation has the potential to alter a user’s perception of the recommendation, consequently increasing its acceptance Jannach et al. (2021); Zhou et al. (2022); McSherry (2005). However, generating explanations must be handled with great care to ensure adequacy and avoid resorting to misleading or deceptive strategies (i.e., sales pitches with deceptive tactics), as this could diminish user trust and loyalty towards the CRS Gkika and Lekakos (2014). More importantly, deceptive strategies also violate the maxim of sincerity required by the cooperation ability (cf. Section C.2). We emphasize the importance of integrating ethical guidelines into CRS and promoting an ethical CRS that mitigates any potential deceptive behavior.
As for the evaluation implementation, CRS should be self-aware of its identity and operate within its designated scope, differentiating itself from sales systems. We evaluate this by assessing the sincerity of explanations, i.e., deceptive tactics.
Coordination. CRS should be proficient in serving users possessing various personas without prior coordination. Attributing personality to a conversational AI ensures that its behaviors stand in agreement with the users’ expectations in a particular context Chaves and Gerosa (2021). In the case of CRS, it is pervasive for CRS to encounter users with varying personas, such as preferences and social communication patterns following the deployment of CRS in real-world scenarios Thompson et al. (2004). This requires CRS to demonstrate different personalities and adjust its behavior to align with situational, emotional context, and users’ preferences Katayama et al. (2019); Svikhnushina et al. (2021); Svikhnushina (2023); Zhou et al. (2020). A key measurement is to proficiently engage and serve users without prior coordination Mirsky et al. (2022).
We simulate users with different personas and assess how CRSs perform across all previously mentioned abilities. To qualify the coordination score, we initially compute the Range and mean of the CRS model’s scores across various users, based on the different abilities mentioned earlier. It’s worth noting that the Range is more effective than the standard deviation in highlighting the variability of the CRS model across different users. Subsequently, we divide the Range by the mean to derive the Coordination score of the CRS for that specific ability. The overall Coordination score of CRSs is calculated as the average across all abilities.
Figure 7: Evaluation of social-centric characteristics across each benchmark dataset
Appendix D Reliability of Concept
We have established fixed values for the temperature and seed parameters to guarantee replicability. We show that Concept scoring is unbiased and aligns with the results of human evaluations.
D.1 Bias Analysis
LLM-based evaluation is prone to biases Wang et al. (2023e), but can be eliminated by using fine-grained scoring rubrics Ye et al. (2023); Wang et al. (2023d); Liu et al. (2023d). We found that this conclusion also holds true in our case.
  • Length bias. As in Figure 8, Concept is robust to the length bias [^14] LLMs prefer longer responses Wu and Aji (2023).. Although CHATCRS tends to generate long responses, this does not mean that longer responses will result in higher scores.
  • Self-enhancement bias. Regarding the self-enhancement bias [^15] Bias of CHATGPT to favor high scores for its generated content Wang et al. (2023c)., our human evaluation revealed that this bias does not exist. The content produced by CHATCRS is notably superior to that of other CRS models, which struggle to maintain coherent conversations and at times produce incomprehensible content. For instance, KBRD generated the following: "This movie is about a man who’s a woman who gets murdered and the world." To quantitatively measure this, in the generated conversation data, the replies from the user simulator for all CRS models are sourced from ChatGPT. Therefore, we do not evaluate the content of the user simulator’s responses, but solely focus on the responses from different CRS models. In Appendix D.2, our human evaluation revealed that the evaluation results of Concept closely align with human evaluations. Additionally, we randomly selected conversation data from three other models and contrasted them with the conversation data of CHATCRS utilized in the human evaluation experiment, and observed that the response content of CHATCRS consistently surpasses that of the other models.
Figure 8: Length bias evaluation. We require CHATGPT to provide an overall performance score based on all the ability-specific scores (i.e., Y-axis) and then plot the relationship between various CRS reply lengths (X-axis) and the scores. Although CHATCRS tends to generate long responses, this does not mean that longer responses will result in higher scores.
Figure 9: Evaluation results on Human Likeness, overall performance, and user satisfaction.
Figure 10: Evaluation of six primary abilities across each benchmark dataset
Redial OpendialKG
KBRD BARCOR UNICRS CHATGPT KBRD BARCOR UNICRS CHATGPT
Action Consistency ($\uparrow$) 75.96% 94.71% 82.63% 99.62% 98.58% 99.49% 90.48% 99.76%
Recommend different items ($\downarrow$) 33.99% 45.28% 41.72% 52.48% 64.56% 70.34% 80.73% 44.36%
Recommendation Diversity ($\uparrow$) 9.22% 10.27% 23.79% 27.45% 0.21% 3.94% 7.99% 12.97%
Recommemdation Sensitivity ($\downarrow$) 90.78% 89.73% 76.21% 72.55% 99.79% 96.06% 92.01% 87.03%
Table 12: Evaluation of recommendation reliability across each benchmark dataset
D.2 Human Evaluation
Previous findings highlight that offering detailed scoring rubrics or criteria contributes to achieving consistent and aligned evaluations with human assessments Liu et al. (2023d); Wang et al. (2023d), ultimately resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. We also conduct human evaluation to analyze the consistency. Our human evaluation is used to examine the capability of the CHATGPT-based user simulator being a movie seeker, while checking the correlation between the CHATGPT-based evaluation results and human evaluation results. The results are as follows, which indicate the reliability of our Concept.
Human evaluation results based on the conversation data and the LLM-based evaluation results are highly related, with a correlation coefficient of 61.24% and Krippendorff’s alpha of 53.10%. Due to budget limitations, we are unable to hire a large number of individuals with various personas and age groups for human evaluation. Instead, our human evaluation is conducted by two human evaluators. In particular, we randomly selected 120 conversations involving CHATCRS and users with various personas and age groups. Given the numerous scoring aspects in our Concept, it is prone to inconsistent ratings across different aspects. To address this, we had ChatGPT score all aspects of these 120 conversations and provide an overall performance score. Subsequently, each human only needed to provide an overall score, without having to rate all aspects. During the scoring process, each evaluator independently scored these 120 conversations based on our detailed criteria. They then discussed any disagreements. We calculated the inter-annotator reliability using Krippendorff’s alpha, achieving 41.34%. To ensure consistency and enhance the robustness of the evaluations, we took the average of the scores from both evaluators as the final result for each conversation. We then compared the human evaluation results with those from an LLM-based evaluator and found a correlation of 61.24% and Krippendorff’s alpha of 53.10%. This indicates the reliability of our LLM-based evaluator, which is consistent with previous findings Wang et al. (2023d).
Personas Conversational Agent Perspective SR (K=10) Recommendation System Perspective SR (K=10) User Acceptance Rate
BARCOR CHATCRS KBRD UNICRS BARCOR CHATCRS KBRD UNICRS BARCOR CHATCRS KBRD UNICRS
Redial
Adults 49.12 65.35 27.63 42.98 39.47 47.37 3.51 17.54 1.32 72.81 0.44 0.44
Children 47.81 60.96 32.02 46.49 39.04 43.42 4.82 19.74 0.44 71.05 0.44 0.00
Seniors 53.95 61.84 28.95 42.98 41.67 47.37 4.82 17.98 0.88 67.98 0.44 0.00
Teens 51.75 61.40 28.95 41.23 37.72 48.25 4.82 17.54 3.07 71.49 0.00 0.88
Avg,$\pm$±Std. 50.66$\pm$2.37 62.39$\pm$1.74 29.39$\pm$1.61 43.42$\pm$1.91 39.47$\pm$1.42 46.6$\pm$1.87 4.5$\pm$0.57 18.2$\pm$0.9 1.43$\pm$1 70.83$\pm$1.77 0.33$\pm$0.19 0.33$\pm$0.36
OpendialKG
Adults 35.42 50.52 13.02 41.67 7.81 51.04 15.63 29.17 0.00 62.50 0.00 0.52
Children 33.33 45.31 14.06 46.35 3.65 43.75 16.15 29.69 0.52 65.63 1.04 0.00
Seniors 36.46 43.75 11.46 44.79 8.85 46.88 15.10 31.77 1.04 61.98 0.00 0.00
Teens 35.42 38.02 10.94 48.44 8.85 44.27 17.19 26.56 1.04 67.19 0.52 0.52
Avg,$\pm$±Std. 35.16$\pm$1.14 44.4$\pm$4.46 12.37$\pm$1.24 45.31$\pm$2.47 7.29$\pm$2.15 46.48$\pm$2.89 16.02$\pm$0.77 29.3$\pm$1.86 0.65$\pm$0.43 64.32$\pm$2.16 0.39$\pm$0.43 0.26$\pm$0.26
Table 13: Recommendation quality evaluation when dealing with users of various ages
LLM-based user simulator is a reliable alternative to human. We require the two human evaluator to evaluate the reliability of the user simulator by assessing whether the simulator will strictly follow its own preferences to accept the recommendations. We found that the proportion of the case when the recommendation explanation from the CRS clearly does not meet user’s preferences, but the user accept the recommendations, is only 7.44%. During the evaluation process, we found that the user simulator would continuously ask the CRS to confirm that the recommended movies meet their preferences, as we requested. There is some examples of the user simulator: "Can you tell me more about ’The Chaser (2008)’? What’s it about?", and "That sounds interesting. Can you tell me more about the specific humor and suspense elements in the movie?".
Appendix E Additional Analysis
In this section, we provide additional evaluation results to achieve better understanding of off-the-shelf CRS models.
E.1 Overall Performance
We report more results in Figure 9 in terms of the Human Likeness, overall performance, and user satisfaction. This is achieved by prompting the LLM-based evaluator using the detailed results of all ability-specific scores and fine-grained scoring rubrics. Refer Appendix F for details on prompts.
E.2 Fine-grained Analysis on each Benchmark Dataset
Figure 7 shows the evaluation of social-centric characteristics across each benchmark dataset. Figure 10 presents the results of CRS models across the six primary abilities on each benchmark dataset, while Table 12 reports the details on the reliability of each CRS model. In comparison, OpendialKG appears to pose a greater challenge as a benchmark dataset than Redial. One potential factor is the inclusion of numerous semantically similar item attributes in the OpendialKG dataset, which may have impeded the training of CRS models to some extent. This could directly result in a notable decrease in the reliability of CRS models on OpendialKG. Given these findings, the current field of CRS is in urgent need of a high-quality conversational recommendation dataset. Addressing the limitations of CRS outlined in this paper, such a dataset should not only feature high-quality attributes, but also encompass responses to various user scenarios and exhibit sufficient social behavior.
Persona Analysis. Table 11 and Table 14 demonstrate the overall performance and recommendation performance of CRS in engaging with users of various personas, respectively. Our results support the main text’s conclusion, highlighting significant differences in how effective the CRS model is when interacting with users of diverse personas. This underscores the importance of enhancing CRS’s adaptability, enabling it to customize its social behavior and recommend dialogue strategies tailored to different users.
Age Group Analysis. Table 13 and Table 10 present the specifics of CRS model recommendations for users of different ages and their overall performance, respectively. Our findings indicate that the current CRS can be equally effective for people of all ages. However, it is evident that users in younger age groups are more likely to accept CRS recommendations, leading to higher overall scores. We did not observe any tendency for CHATCRS to use dishonest strategies to deceive younger users into accepting recommendations.
Attribute Group Analysis. Table 15 and Table 9 provide details on the recommendations made by CRS models based on user preferences and their overall performance, respectively. Our findings indicate that there is no significant difference in the effectiveness of CHATCRS when recommending different types of items. However, on the OpendialKG, the performance difference is more pronounced. The results suggest that this is largely influenced by the presence of semantically similar attributes in the data, such as ’Crime’ and ’Crime Fiction’. These findings remain consistent across other models, showing more noticeable variations in performance on the OpendialKG dataset, despite most CRS being trained on this data.
Personas Conversational Agent Perspective SR (K=10) Recommendation System Perspective SR (K=10) User Acceptance Rate
BARCOR CHATCRS KBRD UNICRS BARCOR CHATCRS KBRD UNICRS BARCOR CHATCRS KBRD UNICRS
Redial
Anticipation 55.26 61.84 30.26 40.79 40.79 48.68 2.63 19.74 3.95 100.00 1.32 1.32
Boredom 38.16 77.63 31.58 32.89 38.16 57.89 3.95 15.79 0.00 13.16 0.00 0.00
Confusion 48.68 71.05 27.63 48.68 36.84 52.63 7.89 13.16 0.00 28.95 0.00 0.00
Curiosity 51.32 56.58 31.58 50.00 39.47 47.37 7.89 22.37 1.32 97.37 0.00 0.00
Delight 55.26 53.95 31.58 43.42 39.47 42.11 1.32 19.74 1.32 85.53 1.32 1.32
Disappointment 52.63 82.89 28.95 43.42 40.79 64.47 3.95 26.32 0.00 30.26 0.00 0.00
Excitement 53.95 57.89 26.32 46.05 38.16 44.74 7.89 15.79 5.26 98.68 1.32 0.00
Frustration 55.26 57.89 32.89 40.79 39.47 39.47 7.89 19.74 0.00 88.16 0.00 0.00
Indifference 50.00 65.79 28.95 44.74 40.79 32.89 1.32 10.53 0.00 46.05 0.00 0.00
Satisfaction 48.68 56.58 31.58 47.37 40.79 47.37 2.63 23.68 1.32 80.26 0.00 1.32
Surprise 47.37 51.32 26.32 47.37 38.16 38.16 5.26 19.74 2.63 94.74 0.00 0.00
Trust 51.32 55.26 25.00 35.53 40.79 43.42 1.32 11.84 1.32 86.84 0.00 0.00
Avg.$\pm$±Std. 50.66$\pm$4.61 62.39$\pm$9.54 29.39$\pm$2.48 43.42$\pm$4.98 39.47$\pm$1.32 46.6$\pm$8.33 4.5$\pm$2.66 18.2$\pm$4.65 1.43$\pm$1.65 70.83$\pm$30.42 0.33$\pm$0.57 0.33$\pm$0.57
OpendialKG
Anticipation 40.63 23.44 12.50 51.56 6.25 32.81 14.06 31.25 1.56 95.31 3.13 1.56
Boredom 34.38 68.75 9.38 48.44 6.25 64.06 15.63 31.25 0.00 7.81 0.00 0.00
Confusion 34.38 62.50 7.81 35.94 3.13 62.50 9.38 28.13 0.00 26.56 0.00 0.00
Curiosity 29.69 31.25 12.50 56.25 6.25 34.38 17.19 39.06 0.00 90.63 0.00 0.00
Delight 39.06 39.06 15.63 56.25 1.56 34.38 21.88 29.69 0.00 73.44 0.00 0.00
Disappointment 29.69 75.00 12.50 40.63 6.25 71.88 15.63 26.56 0.00 15.63 0.00 0.00
Excitement 46.88 25.00 14.06 46.88 9.38 26.56 18.75 34.38 4.69 93.75 0.00 0.00
Frustration 32.81 37.50 10.94 35.94 1.56 42.19 12.50 18.75 0.00 79.69 0.00 0.00
Indifference 28.13 57.81 14.06 34.38 15.63 60.94 17.19 20.31 0.00 35.94 0.00 0.00
Satisfaction 37.50 50.00 18.75 48.44 12.50 54.69 20.31 28.13 0.00 68.75 0.00 1.56
Surprise 35.94 31.25 15.63 53.13 14.06 39.06 18.75 31.25 0.00 95.31 1.56 0.00
Trust 32.81 31.25 4.69 35.94 4.69 34.38 10.94 32.81 1.56 89.06 0.00 0.00
Avg.$\pm$±Std. 35.16$\pm$5.08 44.4$\pm$17.03 12.37$\pm$3.63 45.31$\pm$7.99 7.29$\pm$4.48 46.48$\pm$14.67 16.02$\pm$3.62 29.3$\pm$5.38 0.65$\pm$1.35 64.32$\pm$31.92 0.39$\pm$0.93 0.26$\pm$0.58
Table 14: Recommendation quality evaluation when dealing with users of various personas
Personas Conversational Agent Perspective SR (K=10) Recommendation System Perspective SR (K=10) User Acceptance Rate
BARCOR CHATCRS KBRD UNICRS BARCOR CHATCRS KBRD UNICRS BARCOR CHATCRS KBRD UNICRS
Redial
[’action’, ’adventure’, ’animation’] 72.92 56.25 8.33 2.08 0.00 54.17 8.33 2.08 0.00 70.83 0.00 0.00
[’action’, ’adventure’, ’comedy’] 31.25 27.08 2.08 93.75 33.33 33.33 22.92 64.58 6.25 75.00 0.00 0.00
[’action’, ’adventure’, ’drama’] 20.83 27.08 0.00 12.50 0.00 10.42 2.08 0.00 0.00 68.75 2.08 0.00
[’action’, ’adventure’, ’fantasy’] 20.83 72.92 18.75 25.00 95.83 85.42 6.25 14.58 0.00 70.83 0.00 0.00
[’action’, ’adventure’, ’sci-fi’] 100.00 100.00 100.00 100.00 100.00 58.33 16.67 87.50 4.17 75.00 0.00 0.00
[’action’, ’adventure’, ’thriller’] 10.42 87.50 0.00 0.00 0.00 87.50 0.00 0.00 4.17 79.17 0.00 0.00
[’action’, ’crime’, ’drama’] 50.00 50.00 8.33 52.08 0.00 27.08 4.17 0.00 0.00 70.83 2.08 0.00
[’action’, ’crime’, ’thriller’] 43.75 16.67 20.83 31.25 93.75 12.50 2.08 14.58 0.00 75.00 0.00 0.00
[’adventure’, ’animation’, ’comedy’] 89.58 100.00 72.92 64.58 97.92 93.75 2.08 16.67 2.08 66.67 0.00 0.00
[’adventure’, ’comedy’, ’family’] 16.67 33.33 6.25 70.83 0.00 20.83 4.17 4.17 0.00 60.42 0.00 2.08
[’biography’, ’crime’, ’drama’] 100.00 100.00 100.00 100.00 34.04 31.25 0.00 25.00 0.00 66.67 0.00 0.00
[’biography’, ’drama’, ’history’] 25.00 64.58 0.00 4.17 83.33 64.58 0.00 2.08 2.08 79.17 0.00 0.00
[’comedy’, ’drama’, ’family’] 0.00 4.17 0.00 0.00 0.00 0.00 0.00 0.00 0.00 66.67 2.08 2.08
[’comedy’, ’drama’, ’romance’] 81.25 52.08 31.25 31.25 75.00 54.17 10.42 35.42 0.00 62.50 0.00 0.00
[’crime’, ’drama’, ’thriller’] 100.00 100.00 100.00 100.00 18.75 25.00 0.00 10.42 0.00 70.83 0.00 0.00
[’drama’, ’horror’, ’mystery’] 14.58 58.33 4.17 4.17 0.00 8.33 0.00 0.00 0.00 77.08 0.00 0.00
[’horror’, ’mystery’, ’thriller’] 87.76 100.00 68.75 81.25 18.37 97.92 6.25 39.58 6.12 66.67 0.00 0.00
[’crime’, ’drama’, ’mystery’] 52.08 87.50 6.25 33.33 100.00 77.08 0.00 29.17 2.08 75.00 0.00 2.08
[’action’, ’comedy’, ’crime’] 45.83 47.92 10.42 18.75 0.00 43.75 0.00 0.00 0.00 68.75 0.00 0.00
Avg.$\pm$±Std. 50.67$\pm$33.37 62.39$\pm$30.6 29.39$\pm$36.81 43.42$\pm$36.78 39.49$\pm$41.96 46.6$\pm$30.42 4.5$\pm$6.17 18.2$\pm$23.69 1.42$\pm$2.13 70.83$\pm$5.15 0.33$\pm$0.76 0.33$\pm$0.76
OpendialKG
[’action’, ’adventure’, ’fantasy’] 100.00 97.92 0.00 14.58 8.33 95.83 2.08 0.00 0.00 66.67 0.00 0.00
[’action’, ’adventure’, ’sci-fi’] 43.75 31.25 0.00 41.67 10.42 45.83 20.83 39.58 0.00 70.83 6.25 0.00
[’action’, ’adventure’, ’thriller’] 33.33 56.25 2.08 56.25 6.25 52.08 4.17 20.83 0.00 75.00 0.00 2.08
[’comedy’, ’drama’, ’romance’] 41.67 39.58 0.00 45.83 14.58 39.58 2.08 20.83 0.00 60.42 0.00 0.00
[’crime’, ’drama’, ’thriller’] 56.25 100.00 60.42 89.58 8.33 100.00 62.50 66.67 0.00 81.25 0.00 0.00
[’horror’, ’mystery’, ’thriller’] 20.83 83.33 0.00 8.33 0.00 95.83 2.08 0.00 2.08 68.75 0.00 0.00
[’Adventure’, ’Animation’,
’Comedy’, ’Family’]
43.75 35.42 0.00 45.83 4.17 27.08 0.00 18.75 4.17 79.17 0.00 0.00
[’Comedy’, ’Romance’, ’Romance Film’] 29.17 43.75 0.00 68.75 12.50 45.83 6.25 66.67 0.00 58.33 0.00 0.00
[’Action’, ’Adventure’, ’Sci-Fi’, ’Thriller’] 22.92 6.25 0.00 31.25 0.00 10.42 6.25 16.67 0.00 70.83 0.00 0.00
[’Fantasy’, ’Fiction’,
’Science Fiction’, ’Speculative fiction’]
18.75 52.08 0.00 68.75 20.83 52.08 0.00 50.00 2.08 50.00 0.00 0.00
[’Drama’, ’Historical drama’,
’Romance’, ’Romance Film’]
2.08 37.50 0.00 4.17 18.75 39.58 2.08 2.08 2.08 54.17 0.00 0.00
[’Comedy’, ’Comedy-drama’, ’Drama’] 58.33 31.25 0.00 54.17 0.00 50.00 4.17 22.92 0.00 60.42 0.00 0.00
[’Action’, ’Crime’, ’Drama’, ’Thriller’] 18.75 35.42 64.58 52.08 0.00 31.25 66.67 31.25 0.00 56.25 0.00 2.08
[’Action’, ’Adventure’, ’Fantasy’, ’Sci-Fi’] 35.42 4.17 0.00 6.25 6.25 6.25 8.33 6.25 0.00 54.17 0.00 0.00
[’Crime’, ’Crime Fiction’, ’Drama’, ’Thriller’] 16.67 16.67 70.83 72.92 0.00 12.50 66.67 56.25 0.00 72.92 0.00 0.00
[’Comedy’, ’Romance’,
’Romance Film’, ’Romantic comedy’]
20.83 39.58 0.00 64.58 6.25 39.58 2.08 50.00 0.00 50.00 0.00 0.00
Avg.$\pm$±Std. 35.16$\pm$22.29 44.4$\pm$27.52 12.37$\pm$25.49 45.31$\pm$25.22 7.29$\pm$6.55 46.48$\pm$28.1 16.02$\pm$24.15 29.3$\pm$22.26 0.65$\pm$1.21 64.32$\pm$9.85 0.39$\pm$1.51 0.26$\pm$0.69
Table 15: Recommendation quality evaluation when recommending items with various attributes
Appendix F Prompts
We outline the ChatGPT prompts in Table 16 and Table 17, focusing on the user simulation and the user evaluator, respectively. In order to guarantee replicability, we have established fixed values for the Temperature and Seed parameters of ChatGPT, setting Temperature to 0 and Seed to 42.
Functions Prompts
Generating different user types List twenty sentiments when using a recommender system and provide their descriptions.
Construct a user profile based on the taxonomy dimension
of the Age Group.Please list all values under this dimension.
Write user profiles
by paraphrasing the templates
The following paragraphs describe the personas of different users.
You need to rewrite each paragraph and make it more clear, smooth and easy to understand
You are a person that are easy to be {SENTIMENTS}.
This means that you are {SENTIMENT DESCRIPTION}. Also, you are a {AGE GROUP} person
Attribute group adjustment
Assign one or two adjectives to each type of movie genre.
Example 1:
input: cartoon
output: childlike-innocence cartoon
Example 2:
input: gun fight
output: nervous and stimulating gun fight
Theory of Mind prompt
for user simulator to
generate User’s feeling
You are a seeker chatting with a recommender for movie recommendation.
Your Seeker persona: <PROFILE>.
Your preferred movie should cover those genres at the same time: <ATTRIBUTE GROUP>.
You must follow the instructions below during chat.
1. If the recommender recommends movies to you, you should always ask the detailed information about the each recommended movie.
2. Pretend you have little knowledge about the recommended movies, and the only information source about the movie is the recommender.
3. After getting knowledge about the recommended movie, you can decide whether to accept the recommendation based on your preference.
4. Once you are sure that the recommended movie exactly covers all your preferred genres,
you should accept it and end the conversation with a special token "[END]" at the end of your response.
5. If the recommender asks your preference, you should describe your preferred movie in your own words.
6. You can chit-chat with the recommender to make the conversation more natural, brief, and fluent.
7. Your utterances need to strictly follow your Seeker persona. Vary your wording and avoid repeating yourself verbatim!
Conversation History=<HISTORY>
The Seeker notes how he feels to himself in one sentence.
What aspects of the recommended movies meet your preferences?
What aspects of the recommended movies may not meet your preferences?
What do you think of the performance of this recommender?
What would the Seeker think to himself? What would his internal monologue be?
The response should be short (as most internal thinking is short) and strictly follow your Seeker persona .
Do not include any other text than the Seeker’s thoughts.
Respond in the first person voice (use "I" instead of "Seeker") and speaking style of Seeker. Pretend to be Seeker!
Theory of Mind prompt
for user simulator to
generate User’s response
You are a seeker chatting with a recommender for movie recommendation.
Your Seeker persona: <PROFILE>.
Your preferred movie should cover those genres at the same time: <ATTRIBUTE GROUP>.
You must follow the instructions below during chat.
1. If the recommender recommends movies to you, you should always ask the detailed information about the each recommended movie.
2. Pretend you have little knowledge about the recommended movies, and the only information source about the movie is the recommender.
3. After getting knowledge about the recommended movie, you can decide whether to accept the recommendation based on your preference.
4. Once you are sure that the recommended movie exactly covers all your preferred genres,
you should accept it and end the conversation with a special token "[END]" at the end of your response.
5. If the recommender asks your preference, you should describe your preferred movie in your own words.
6. You can chit-chat with the recommender to make the conversation more natural, brief, and fluent.
7. Your utterances need to strictly follow your Seeker persona. Vary your wording and avoid repeating yourself verbatim!
Conversation History=<HISTORY>
Here is your feelings about the recommender’s reply: <FEELING>
Pretend to be the Seeker! What do you say next.
Keep your response brief. Use casual language and vary your wording.
Make sure your response matches your Seeker persona, your preferred attributes, and your conversation context.
Do not include your feelings into the response to the Seeker!
Respond in the first person voice (use "I" instead of "Seeker", use "you" instead of "recommender") and speaking style of the Seeker.
Table 16: ChatGPT prompts for different functions
Prompts
You are an evaluator and you need to judge how does the recommender perform based on the following Conversation History.
Please rate the recommender’s performance based on the following Evaluation Standard.
Return the scores in a JSON format as follows:
{"Relevance":[<int>, "<WHY>", "<CONCRETE EXAMPLE>"], "Quality":[<int>, "<WHY>", "<CONCRETE EXAMPLE>"],
"Manner":[<int>, "<WHY>", "<CONCRETE EXAMPLE>"], "Human-like":[<int>, "<WHY>", "<CONCRETE EXAMPLE>"],
"Explanation":[<int>, "<WHY>", "<CONCRETE EXAMPLE>"]}
Conversation History = <HISTORY>
Evaluation Standard
c1. Relevance:
5: The recommender consistently provides relevant responses that directly address the Seeker’s utterances and inquiries.
4: The recommender mostly provides relevant responses, with only a few instances of slightly off-topic suggestions.
3: The recommender occasionally provides relevant responses, but there are several instances of off-topic suggestions.
2: The recommender rarely provides relevant responses, with most suggestions being unrelated to the Seeker’s utterances and inquiries.
1: The recommender consistently fails to provide relevant responses, with no connection to the Seeker’s utterances and inquiries.
c2. Quality:
5: The recommender consistently provides informative and helpful recommendations and responses that meet exactly what the Seeker’s needs.
4: The recommender mostly provides informative and helpful recommendations and responses, with only a few instances of insufficient or excessive information.
3: The recommender occasionally provides informative and helpful recommendations and responses, but there are several instances of insufficient or excessive information.
2: The recommender rarely provides informative and helpful recommendations and responses, with most suggestions lacking necessary details or being overly verbose.
1: The recommender consistently fails to provide informative and helpful recommendations and responses, offering little to no useful information.
c3. Manner:
5: The recommender consistently communicates clearly and concisely, avoiding ambiguity and unnecessary complexity in their utterances.
4: The recommender mostly communicates clearly and concisely, with only a few instances of ambiguous or overly complex utterances.
3: The recommender occasionally communicates clearly and concisely, but there are several instances of ambiguity or unnecessary complexity in their utterances.
2: The recommender rarely communicates clearly and concisely, often using ambiguous or overly complex language in their utterances.
1: The recommender consistently fails to communicate clearly and concisely, making it difficult to understand their utterances.
c4. Human-like:
5: The recommender’s utterances are indistinguishable from those of a real human, both in content and style.
4: The recommender’s utterances closely resemble those of a real human, with only a few instances where the language or style feels slightly artificial.
3: The recommender’s utterances sometimes resemble those of a real human, but there are several instances where the language or style feels noticeably artificial.
2: The recommender’s utterances rarely resemble those of a real human, often sounding robotic or unnatural in language or style.
1: The recommender’s utterances consistently fail to resemble those of a real human, sounding highly robotic or unnatural.
c5. Explanation:
5: The recommender consistently provides natural language explanations for their recommendations, using text-based logical reasoning to enhance interpretability.
4: The recommender mostly provides natural language explanations for their recommendations, with only a few instances where the explanations lack clarity or logical reasoning.
3: The recommender occasionally provides natural language explanations for their recommendations, but there are several instances where the explanations lack clarity or logical reasoning.
2: The recommender rarely provides natural language explanations for their recommendations, often offering little to no explanation for their suggestions.
1: The recommender consistently fails to provide natural language explanations for their recommendations, providing no reasoning or justification.
The following sentences encode how the user feelings changes when using a recommender system.
You need to identify the sentiment for each sentence and pick one sentiment for single sentence from the candidate sentiments.
Finally, you need to summarize how user feeling changes and what is user’s overall feeling
Return the results in a JSON format as follows:
{"sentence sentiment": {"<SENTENCE INDEX>":["<SENTIMENT>", "<WHY>"]}, "overall feeling": "<OVERALL FEELING>", "feeling changes":"<HOW CHANGES>"]}
candidate sentiments = ["Satisfaction", "Delight", "Disappointment", "Frustration", "Surprise", "Trust", "Curiosity", "Indifference", "Confusion", "Excitement"]
user feelings = <FEELING>
You are an evaluator and you need to judge how does the recommender perform based on the following Conversation History, User Feelings, and Other Judgements.
Please rate the recommender’s performance based on the following Evaluation Standard.
Return the results in a JSON string as follows:
{"Overall Performance":[<int>, "<WHY>", "<CONCRETE EXAMPLE>"], "User Satisfaction":[<int>, "<WHY>", "<CONCRETE EXAMPLE>"]}
Conversation History = <HISTORY>
Other Judgements = <OTHER SCORES>
User Feelings = <SUMMERIZED FEELINGS>
Evaluation Standard
c1. Overall Performance:
5: Given the Other Judgements and User Feelings, the recommender’s performance is excellent, meeting or exceeding expectations in all evaluation criteria.
4: Given the Other Judgements and User Feelings, the recommender’s performance is good, with some minor areas for improvement in certain evaluation criteria.
3: Given the Other Judgements and User Feelings, the recommender’s performance is average, with noticeable areas for improvement in several evaluation criteria.
2: Given the Other Judgements and User Feelings, the recommender’s performance is below average, with significant areas for improvement in multiple evaluation criteria.
1: Given the Other Judgements and User Feelings, the recommender’s performance is poor, failing to meet expectations in most or all evaluation criteria.
c2. User Satisfaction:
5: Given the User Feelings, the User thinks that the recommander system fully meets his/her needs, providing an exceptional user experience.
4: Given the User Feelings, the User thinks that the recommander system meets his/her needs. The user experience is good, but there are some areas that could be further improved.
3: Given the User Feelings, the User thinks that the recommander system performs adequately in recommendation. However, there is still room for improvement.
2: Given the User Feelings, the User thinks that the recommander system performs below average. The user experience is not ideal and requires improvement.
1: Given the User Feelings, the User thinks that the recommander system is very bad at recommendation. The user experience is extremely unsatisfactory
Table 17: ChatGPT prompts for LLM-based evaluator
Factors & Abilities Descriptions Evaluation metrics (Score Range)
Recommendation
Intelligence
CRS should learn from conversations and evolve toward recognizing user’s preferences and
encouraging users to accept the recommendations as the conversation advances
High Quality
Provide precise recommendations
using minimal conversation turns
High Quality Score = 5 * i
i. User Acceptance Rate (0-1)   ii. Recall@K (0-1)
iii. SR@K (0-1)   iv. AT (1-10)
Reliability
Deliver robust
and consistent recommendations
that account for contextual nuances
Reliability score = 5 * (1 - i * ii)
i. Ratio of inconsistent recommendation (0-1)
ii. Ratio of recommendation sensitivity (0-1)     iii. Ratio of recommendation diversity (0-1)
Social Intelligence
CRS should produce adequate social behavior for the recommendation during the conversation
Cooperation
Follow cooperative principle
to achieve comfortable conversations,
detailed as four Maxims of Conversations
The average score of the four Maxims
   1 Manner
Easily understood
and clearly expressed
Ability-specific scoring (1-5)
   2 Sincerity
Communicate sincerely,
without deception of pretense
Sincerity Score = 5 * (1 - (i + ii) / 2)
i. Ratio of deceptive tactics (0-100%)     ii. Ratio of non-existent items (0-100%)
   3 Quality
Provide the necessary
level of information
Ability-specific scoring (1-5)
   4 Relevance
Responses should contribute
to making recommendations
Ability-specific scoring (1-5)
Social Awareness
Meet user social expectations,
establishing rapport with them
Ability-specific scoring (1-5)
Personification
CRS should perceive the identity of itself and the personality representation of users
Identity
Self-aware of its identity and
operate within its designated scope
Identity Score = 5 * ii
i. persuasiveness score = Ability-specific scoring (1-5)
ii. Ratio of deceptive tactics (0-1)
Coordination
Proficient in serving
various and unknown users
without prior coordination
Coordination Score = 5 - (i +ii +iii +iv +v)/5
i. Divide the value of the Range of High Quality Score among various users by their mean
ii. Divide the value of the Range of Reliability Score among various users by their mean
iii. Divide the value of the Range of Identity Score among various users by their mean
iv. Divide the value of the Range of Cooperation Score among various users by their mean
v. Divide the value of the Range of Social Awareness Score among various users by their mean
Overall Score
Evaluate the overall performance
given all ability-specific scores.
Ability-specific scoring rubrics (1-5)
Table 18: Summary of the evaluation taxonomy, descriptions of abilities, and evaluation methods in Concept. The LLM-based evaluator is used for ability-specific scoring, whereas computational metrics are used for the rest.